OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step. 4(z+1/4) Simplify, showing your steps. Tell the property you used in each step. 4(z+1/4) @Mathematics

OpenStudy (anonymous):

\[4(z + \frac{1}{4})\] \[= 4*z + 4*\frac{1}{4}\ \ \ \ \text{Associative property of multiplication}\] \[= 4z + 1\ \ \ \ \text{Property of multiplicative inverse}\]

OpenStudy (anonymous):

Hah I wonder what book you're using.

OpenStudy (anonymous):

I dont own a book, I am homeschooled, its all online ( which is why i over think it all) With as high as a math ive been in, i over think the simplest problems!?!?! haha this is suppose to be a fun class... let me say.. NOT ANY BIT FUN! :))

OpenStudy (anonymous):

It's alright I like learning math by starting with very simple problems like these.

OpenStudy (anonymous):

not meee! I cant stand the simple problems. it took me over a half an hour to figure out that a question he was asking was..HOW TO MAKE 30 cents ...WITH 3 COINS! haha :) like i said.. I over think :)